OpenStudy (anonymous):

I hate fractions, how do i solve this: 4-9/25m^2

OpenStudy (anonymous):

Its a binomial

OpenStudy (ash2326):

4- 9/25 *m^2=0 Is this your question?

OpenStudy (anonymous):

theres no 0

OpenStudy (anonymous):

\[4-\frac{9}{25m^2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):



OpenStudy (anonymous):

hey how did you make your drawing so big?

OpenStudy (ash2326):

ok, then you want to factor it 4-9/25 m^2 \[2^2-(3/5* m)^2\] a^2-b^2=(a+b)(a-b) (2-3/5 m)*(2+3/5 m)

OpenStudy (anonymous):

Im not sure

OpenStudy (ash2326):

Tomas you can drag that window to increase its size??

OpenStudy (ash2326):

*no question marks

OpenStudy (anonymous):

They dont have any common factors do they?

OpenStudy (anonymous):

how? i can't manage to do it

OpenStudy (ash2326):

no, they don't

OpenStudy (anonymous):

he used this rule: \[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

ok i got it, thanks :)